OpenStudy (anonymous):

What is the value of 6!? A. 36 B. 720 C. 5040 D. 46,656

OpenStudy (rational):

\(6! = 6\times 5 \times 4 \times 3 \times 2 \times 1\)

mathslover (mathslover):

Since, n! = n * (n-1) * (n-2) * (n-3) * .... * 3 * 2 * 1 So, 6! = 6 * (6-1) * (6-2) * (6-3) * (6-4) * (6-5) = 6 * 5 * 4 * 3 * 2 * 1

mathslover (mathslover):

Calculate it by yourself or by using calculator.

OpenStudy (anonymous):

What is the value of 9P5 ? A. 45 B. 126 C. 3024 D. 15,120

mathslover (mathslover):

Please post it as another question.